Solve using the elimination method. Use a graphing calculator to check your answers.
8x-5y=77
-6x-3y=-51

The answer they got was (9,-1) How did they get this?

Answer by TimothyLamb(4379)

8x-5y=77
-6x-3y=-51
---
put the system of linear equations into standard form
---
8x-5y=77
-6x-3y=-51
---
copy and paste the above standard form linear equations in to this solver:
https://sooeet.com/math/system-of-linear-equations-solver.php
---
solution:
x = 9
y = -1
---
now by elimination:
---
3(8x - 5y = 77)
4(-6x - 3y = -51)
---
24x - 15y = 231
-24x - 12y = -204
---
add the two equations:
---
-27y = 27
y = 27/-27
y = -1
---
8x - 5y = 77
8x - 5(-1) = 77
8x + 5 = 77
8x = 77 - 5
x = 72/8
x = 9
---
solution by elimination:
x = 9
y = -1
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php
---